OpenStudy (anonymous):

Solve by factoring: 3x^2=6x Please show work, thank you.

OpenStudy (anonymous):

\[3x^2 = 6x\]\[3x^2 - 6x = 0\]\[3x(x - 2) = 0\] this equation is true when x = 0 or when x = 2